OpenStudy (anonymous):

What is the value of the expression shown below? 3 over 5 to the power of 2 + 4 × 3 − 2

OpenStudy (anonymous):

@DecentNabeel @Mertsj

OpenStudy (decentnabeel):

2 + 4 × 3 − 2 2+12-2 14-2 =12 (3/5)^12 =

OpenStudy (anonymous):

the answer would be?

OpenStudy (anonymous):

12?

OpenStudy (decentnabeel):

\[\left(\frac{3}{5}\right)^{12}=\frac{3^{12}}{5^{12}}\quad \left(\mathrm{Decimal:\quad }\:0.00218\right)\]

OpenStudy (anonymous):

12 9/25?

OpenStudy (decentnabeel):

this is new qustion..

OpenStudy (anonymous):

i mean whats the answer to the question

OpenStudy (decentnabeel):

12*9/25 .. you need answer for this question

OpenStudy (anonymous):

whats the answer to the first question

OpenStudy (decentnabeel):

what is the question .. tell me again

OpenStudy (anonymous):

3 over 5 to the power of 2 + 4 × 3 − 2

OpenStudy (decentnabeel):

first solve this one.. 2 + 4 × 3 − 2

OpenStudy (anonymous):

12

OpenStudy (decentnabeel):

gud then 3/5 to the power of 12 then solve

OpenStudy (decentnabeel):

mean.. 12^(3/5) ok

OpenStudy (anonymous):

9.6?

OpenStudy (anonymous):

whats the answer?

OpenStudy (decentnabeel):

\[12^{\frac{3}{5}}=4.44129\] that is the answer

OpenStudy (anonymous):

actually no these are the choices.... 10 and 9 over 25 11 and 2 over 25 12 and 9 over 25 13 and 2 over 25

OpenStudy (anonymous):

I need the answer to the whole question.. i dont have time t6o solve it out.

OpenStudy (anonymous):

*to

OpenStudy (usukidoll):

Due to OpenStudy's Code of Conduct, we can't give out direct answers. We can only guide the user through the process of solving their question.

OpenStudy (anonymous):

ik but solving it is making me even more confused ik the rules

OpenStudy (usukidoll):

But I got to ask... where did you guys get the exponent 12 from when the original question is 3 over 5 to the power of 2 + 4 × 3 − 2 ???

OpenStudy (usukidoll):

wouldn't the problem be something like this? \[(\frac{3}{5})^2+4 \times 3 - 2\]

OpenStudy (usukidoll):

I do apologize the latex isn't working correctly...  is that your question @KAKES1967